d. A resistance strain gage with a gage factor of 2.4 is mounted on a steel beam whose modulus of 

elasticity is 2x106kg/cm2

. The strain gage has an unstrained resistance of 120.0Ω which 

increases to 120.1Ω when the beam is subjected to a stress. Calculate the stress at the point 

where the strain gage is mounted


1
Expert's answer
2021-05-31T11:34:30-0400

Gauge factor is the ratio of relative change in resistance to that of the strain.

Relative change in resistance "\frac{120.1-120}{120}=8.34*10^{-4}"

Gauge factor "=\frac{8.34*10^{-4}}{\epsilon}=2.4"

"Strain, \epsilon = \frac{8.34*10^{-4}}{2.4}=3.5*10^{-4}"